OpenStudy (anonymous):

Let 5 x^2+ 3 x + xy = 1 and y( 1 ) = -7. Find y'( 1 ) by implicit differentiation。 Thx

OpenStudy (anonymous):

\[10x+3+xy'+y=0\]\[y'=\frac{ (-10x-3-y) }{ x }\]\[y'=\frac{ (-10(1)-3+7) }{(1) }\]=-6